OpenStudy (anonymous):

The tangent to the curve y = x3 at (1,1) meets the graph at a second point. Having trouble getting the second point, without just resorting to guessing..

OpenStudy (amistre64):

define the tangent line at the point 1,1

OpenStudy (amistre64):

then equate where the tangent line and the curve are equal

OpenStudy (amistre64):

you should have something like: y=x^3 and y=mx+b therefore the points in common are when: x^3 = mx+b

OpenStudy (anonymous):